OpenStudy (anonymous):

Gabby solve {x^2} - x - 20 = 0 as shown below: \begin{array}{l}{x^2} - x - 20 = 0\\\left( {x - 5} \right)\left( {x + 4} \right) = 0\\x = - 5,4\end{array} Caitlyn says Gabby is wrong and believe this is the solution: \begin{array}{l}{x^2} - x - 20 = 0\\\left( {x - 5} \right)\left( {x + 4} \right) = 0\\x = 5, - 4\end{array} Is Gabby or Caitlyn correct? Explain the error made by the incorrect girl.

OpenStudy (solomonzelman):

Well, first we have to know that when we say `a•b=0` (in general) THEN, either a or b has to be equal to 0. is this information correct?

OpenStudy (anonymous):

by correct you mean the way it was typed ?

OpenStudy (solomonzelman):

I mean that the point I made (as regards to the math part)

OpenStudy (solomonzelman):

that the point I made is correct or not

OpenStudy (anonymous):

{x^2} - x - 20 = 0 this is the problem they are working out

OpenStudy (solomonzelman):

yes, yes, I know.... I am asking you if you are {familiar with/ understand the} the zero product property

OpenStudy (anonymous):

not really

OpenStudy (solomonzelman):

You know if you say `x • y = 0` then for it to be true, either x=0, or y=0

OpenStudy (solomonzelman):

Right?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

Ok, now back to our problem....

OpenStudy (solomonzelman):

he factored the quadratic equation, and now he has: \(\large\color{black}{ \displaystyle (x-5)(x+4)=0 }\)

OpenStudy (solomonzelman):

Then, this "zero product property" (as I showed with `y•x=0`) has to work here too.

OpenStudy (solomonzelman):

That means that either x-5 =0 or that x+4 =0 Correct or not?

OpenStudy (solomonzelman):

Now, for what value of x is x-5=0? And for what value of x is x+4=0?

OpenStudy (anonymous):

it would be 4, and 5 ?

OpenStudy (solomonzelman):

Please in order, and please try again

OpenStudy (anonymous):

it would be -5, 4

OpenStudy (solomonzelman):

x-5=0 x+4=0 (-5)-5=0 (4)+4=0 according to what you said:

OpenStudy (anonymous):

so the property that they did wrong was the zero property

OpenStudy (solomonzelman):

who did wrong?

OpenStudy (anonymous):

Caitlyn and Gabby

OpenStudy (solomonzelman):

One of them is correct. (But only one is correct, the other is wrong)

OpenStudy (triciaal):